OpenStudy (anonymous):

how do you evaluate 28 divided by tree and a half

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Is it 28 divided by three and a half??? if yes, u can do as follows 2 56 28 / (3 1/2) = 28 / (7/2) = 28 x --- = ---- = 8 7 7